Use the substitution method to solve the following system of equation
2x-9y=1
x-4y=9

Answer by algebrahouse.com(1659)

"Use the substitution method to solve the following system of equations"

2x - 9y = 1
x - 4y = 9 ---------> x = 4y + 9 {added 4y to both sides}

2x - 9y = 1 {top equation}
2(4y + 9) - 9y = 1 {substituted 4y + 9, in for x, into top equation}
8y + 18 - 9y = 1 {used distributive property}
-y + 18 = 1 {combined like terms}
-y = -17 {subtracted 18 from both sides}
y = 17 {divided both sides by -1}

x = 4y + 9
x = 4(17) + 9 {substituted 17, in for y, into 4y + 9}
x = 68 + 9 {multiplied 4 by 17}
x = 77 {added 68 and 9}

x = 77 and y = 17
www.algebrahouse.com